# Measuring AI Ability to Complete Long Tasks

### v1

@@ 1 Introduction
In the last five years, frontier AI systems have undergone a dramatic transformation in capabilities, evolving from basic text generation [1] to autonomously executing complex multi-hour machine learning research projects [2]. Sufficiently capable AIs could perform dangerous, highly complex actions like autonomous development of chemical, biological, radiological or nuclear weapons (CBRN) and self-replication and adaptation outside human control [3]. Understanding AI capabilities helps inform the development of safety guardrails as systems become increasingly powerful. In particular, many frontier AI developers have committed to using measures of specific AI capabilities to determine the necessary risk mitigations for their frontier AI systems. [^1] An updated list of frontier AI safety policies can be found at: https://metr.org/faisc. Robust benchmarks that can accurately track and forecast AI capabilities thus form the foundation for responsibd AI governance and risk mitigation.
@@
We propose tracking AI progress over time using the task completion time horizon: the duration of tasks that models can complete at a certain success probability, providing an intuitive measure of real-world capability compared to humans. As models may not reliably complete all tasks of a given length, we operationalize this by measuring the X%-(task completion) time horizon–the length of tasks that models can complete approximately X% of the time.
@@
First, we replicate our methods on SWE-bench Verified (Section 6.3), which includes human difficulty annotations. We find that the exponential trend still holds (Figure 11), albeit with an even shorter doubling time, possibly because SWE-bench Verified difficulty annotations may differentially underestimate how long human contractors take to perform easier SWE-bench tasks.
Second, we score our HCAST and RE-Bench tasks against a list of 16 “messiness” factors that aim to capture some systematic differences between our tasks and hard“real-world” tasks. Some examples include whether the task is resource limited, novel, or involves a dynamic environment (Section D.4). Controlling for task length, we find models perform worse on tasks that have higher messiness scores. Notably, trends in AI agent performance over time are similar for the lower and higher messiness subsets of our tasks (see Figure 9). In particular, we find no evidence of plateaus in performance trends specific to our higher messiness subset.
@@
Our supplementary experiments find little evidence of performance trends being slower on the somewhat more realistic tasks we tested, but do not rule out the possibility that trends are meaningfully slower on the distribution of tasks required to automate software engineer jobs. These experiments were designed to detect large trend differences and constant factor shifts in model performance. We also find evidence that AI agent time horizons can differ by a large constant factor depending on the domain and reference human population.
@@ 3.3.1 Agent scaffolds
Most AI models were evaluated with modular-public—our basic agent scaffold. [^11] Code for this agent scaffold can be found at https://github.com/poking-agents/modular-public. This scaffold provides the model with Python and Bash commands and some very simple context management to keep the input within the context window length of the LM. We used a slightly different scaffold for o1-preview and o1, because they seemed to struggle with tool use, responding to environmental feedback, and generally acting as an agent. These are described further in Appendix B.3. In general, all
@@ 4.1 From raw data to time horizon
Figure 4: Model success rates are negatively correlated with how much time it takes a human to complete the task. (y=0.07x+0.66y=-0.07x+0.66, $R^{2}:0.83$)
@@
In Figure 1, we plot the time horizons of each model against their release date—the date at which the lab first publicly announced the frontier model. [^13] In most cases, the release date is for each frontier AI model is the same model as the one we evaluated. However, GPT-3 (davinci) and GPT-3.5 (code-davinci-002 and text-davinci-002) are closed-source and no longer available through API access, but we use their release dates for the closest available models, which OpenAI advertises as having equivalent performance: GPT-3’s release date for davinci-002, and GPT-3.5’s date for gpt3.5-turbo-instruct. In addition, we linearly regress [^14] Specifically, we perform Ordinary Least Squares regression on $\log(\text{model\_horizon})=\alpha+\beta\cdot\text{release\_date}$. In Appendix D we discuss other curve-fitting methods, and conclude that the fit is not sensitive to various hyperparameters such as regularization, weighting of tasks, or WLS vs. OLS. $\log(\text{time horizon})$ against release date, finding that time horizon has doubled every 212 days with a 95% bootstraped confidence interval 171–249 days (or $\pm 19\%$). Error bars are calculated via 10,000 samples from a three-level hierarchical bootstrap over task families, then tasks, then runs.
@@ 4.2.1 Time horizons at 50% success rate vs 80% success rate
To check whether our choice of 50% success rate affects the long-run trend, we also compute the time horizon at which AI agents succeed at tasks with 80% success rate, shown in Figure 6. The doubling time in 80% time horizon (213 days) is similar to the doubling time of 50% time horizon (212 days), within margin of error. However, there is a large gap between models’ 50% time horizon and 80% time horizon—Claude 3.7 Sonnet has the longest 80%-horizon among models we examined at around 15 minutes, in contrast to its 50%-horizon of 59 minutes. This gap suggests that even models that sometimes succeed on difficult and diverse tasks cannot reliably perform tasks of moderate length.
@@ 6 External validity and robustness
To investigate the applicability of our results to other benchmarks, and to real task distributions, we performed four supplementary experiments. First, we check whether the 2023–2025 trend without the SWAA dataset retrodicts the trend since 2019, and find that the trends agree. Second, we label each of our tasks on 16 “messiness” factors—factors that we expect to (1) be representative of how real-world tasks may systematically differ from our tasks and (2) be relevant to AI agent performance. Third, we calculate AI agent horizon lengths from SWE-bench Verified tasks. We find a similar exponential trend, although with a shorter doubling time. However, we believe this shorter doubling time to be a result of SWE-bench Verified time annotations differentially underestimating the difficulty easier SWE-bench tasks. Finally, we collect and baseline a small set of uncontaminated issues from internal METR repositories. We find that our contracted human baseliners take much longer to complete these tasks than repository maintainers. We also find that AI agent performance is worse than would be predicted maintainer time-to-complete but is consistent with contractor time-to-complete, given the AI agent success curves from HCAST + SWAA + RE-Bench tasks shown in Figure 5.
@@ 6.3 SWE-bench Verified
We found that annotator time estimates differentially underestimate how long our contract baseliners take to complete the easiest SWE-bench verified tasks. As a result, our time horizon estimates for SWE-bench Verified (which use annotator times) are likely to underestimate the time horizon of less capable models, in turn shortening doubling times. For more details see Appendix D.3.
@@ 7.2.2 Future changes in time horizon trends
The main inputs to AI research and development are compute and researcher time. If future AI systems are capable of substituting for human research engineers and/or increasing the compute-efficiency of training, the rate of AI progress will increase. We think it is likely that there will be be substantial AI R&D automation once frontier AI time horizon reaches tens of hours, shortening the time from then until one-month-horizon AI.
@@ 8.1 Measuring and interpreting time horizon
An infinite time horizon does not mean an arbitrarily capable AI, merely the ability to complete tasks that take humans an arbitrarily long length of time. If an artificial general intelligence (AGI) is capable of completing all tasks expert humans can with a success rate of at least X%, its X% time horizon will necessarily be infinite. Therefore, if such systems are ever developed, the long-term trend in time horizon will be faster than exponential, with an asymptote at the date of AGI deployment.
@@ Acknowledgments
The authors thank the following reviewers for feedback on draft versions of this paper. Ryan Greenblatt, Aaron Scher, Romeo Dean, Mike Knoop, Jeff Wu, Steve Newman, Rohit Krishnan, Taren Stinebrickner-Kauffman, JS Denain, Jacob Pfau, Seb Krier, Anton Troynikov, Max Henderson, Ajeya Cotra, Max Nadeau, Tamay Besiroglu, Nate Thomas.
@@ Appendix D More ablations and robustness checks
    Different weightings of task families: $\frac{1}{\sqrt{\text{family size}}}$ vs uniform (included in Figure 12); also $\frac{1}{\text{family size}}$
@@ D.1 Alternative curve fits
  • Sophisticated stochastic models that asymptote to infinity often have large uncertainties; see Roodman [2020], which applied a superexponential diffusion model to economic data. As we discuss in Section 7.2, large uncertainty may be appropriate when constructing a prediction interval for a long-term forecast, but quantitatively modeling the potential impact of the factors we mentioned is out of scope of this paper.
@@ D.3 SWE-bench Verified
SWE-bench Verified contains time estimates created by contractors that split tasks into four buckets: <15<15 min fix, 15 minutes–1 hour, 1 hour–4 hours, or >4>4 hours. These estimates were based on the expected time it would take an “engineer who has had a few hours to familiarize themselves with the codebase” to solve the issue.
We verified annotator time buckets by running seven baselines across 6 SWE-bench Verified tasks. We baselined four tasks in the “<15<15 min fix” bucket, and found that they took our baseliners 8, 26, 67, and 84 minutes respectively. We baselined two tasks in the “1 hour–4 hours” bucket, and both baseliners took between 2–3 hours.
@@
>4>4 hours 480.0 min
Table 8: We convert the SWE-bench Verified time annotations into task estimates, by taking the geometric mean of the time annotation. We caution that this likely underestimates the time each issue takes to resolve by a human baseliner without context—notably, we observe that the geometric mean of baseliner time for four randomly sampled tasks in the “<15<15 minute fix” bucket is 32.9 minutes.
@@
As shown in Table 8, we found that annotator time estimates are likely more inaccurate on the “<<15 min fix” time bucket, as in practice baseliners often take much longer than this to complete tasks. This results in easier tasks’ difficulty being underestimated, meaning that the SWE-bench Verified time horizon estimates likely underestimate true model performance for less capable models, which in turn leads to a steeper exponential trend.
@@ Appendix F Author contributions
Brian Goodrich contributed to the development of the modular-public agent, and created the flock and duet agents. Nikola Jurkovic baselined SWAA and RE-Bench tasks and tested RE-Bench tasks. Seraphina Nix contributed to writing and revising the paper and figures. David Rein managed the development and collection of HCAST tasks and oversaw the human baselining process for those tasks. Lucas Jun Koba Sato led the development of the AI agents and collection of agent results on HCAST tasks. Neev Parikh contributed to running external validity experiments.

### v2

@@ 1 Introduction
In the last five years, frontier AI systems have undergone a dramatic transformation in capabilities, evolving from basic text generation [1] to autonomously executing complex multi-hour machine learning research projects [2]. Sufficiently capable AIs could perform dangerous, highly complex actions like autonomous development of chemical, biological, radiological or nuclear weapons (CBRN) and self-replication and adaptation outside human control [3]. Understanding AI capabilities helps inform the development of safety guardrails as systems become increasingly powerful. In particular, many frontier AI developers have committed to using measures of specific AI capabilities to determine the necessary risk mitigations for their frontier AI systems. [^1] An updated list of frontier AI safety policies can be found at: https://metr.org/faisc. Robust benchmarks that can accurately track and forecast AI capabilities thus form the foundation for responsible AI governance and risk mitigation.
@@
We propose tracking AI progress over time using the task completion time horizon: the duration of tasks that models can complete at a certain success probability, providing an intuitive measure of real-world capability compared to humans. As models may not reliably complete all tasks of a given length, we operationalize this by measuring the X%-(task completion) time horizon–the length of tasks that models can complete approximately Xof the time.
@@
First, we score our HCAST and RE-Bench tasks against a list of 16 “messiness” factors that aim to capture some systematic differences between our tasks and hard“real-world” tasks. Some examples include whether the task is resource limited, novel, or involves a dynamic environment (Section D.4). Controlling for task length, we find models perform worse on tasks that have higher messiness scores. Notably, trends in AI agent performance over time are similar for the lower and higher messiness subsets of our tasks (see Figure 9). In particular, we find no evidence of plateaus in performance trends specific to our higher messiness subset.
Second, we replicate our methods on SWE-bench Verified (Section 6.3), which includes human difficulty annotations. We find that the exponential trend still holds (Figure 11), albeit with an even shorter doubling time, possibly because SWE-bench Verified difficulty annotations, meant to represent high-context maintainer time, may differentially underestimate how long human contractors take to perform easier SWE-bench tasks.
@@
Our supplementary experiments find little evidence of performance trends being slower on the somewhat more realistic tasks we tested, but do not rule out the possibility that trends are meaningfully slower on the distribution of tasks required to automate software engineer jobs. These experiments were designed to detect large trend differences and constant factor shifts in model performance. We also find evidence that AI agent time horizons can differ by a large factor depending on the task domain and reference human population.
@@ 3.3.1 Agent scaffolds
Most AI models were evaluated with modular-public—our basic agent scaffold. [^11] Code for this agent scaffold can be found at https://github.com/poking-agents/modular-public. This scaffold provides the model with Python and Bash commands and some very simple context management to keep the input within the context window length of the LM. We used a slightly different scaffold for o1-preview and o1, because they seemed to struggle with tool use, responding to environmental feedback, and generally acting as an agent. These are described further in Appendix B.3.
@@ 4.1 From raw data to time horizon
Figure 4: Model success rates are negatively correlated with how much time it takes a human to complete the task. ($y=-0.07x+0.66$, $R^{2}:0.83$)
@@
[…]
@@ 4.2.1 Time horizons at 50% success rate vs 80% success rate
To check whether our choice of 50% success rate affects the long-run trend, we also compute the time horizon at which AI agents succeed at tasks with 80% success rate, shown in Figure 6. The doubling time in 80% time horizon (213 days) is similar to the doubling time of 50% time horizon (212 days), within margin of error. However, there is a large gap between models’ 50% time horizon and 80% time horizon—Claude 3.7 Sonnet has the longest 80%-horizon among models we examined at around 15 minutes, in contrast to its 50%-horizon of 59 minutes. This gap suggests that even models that sometimes succeed on difficult and diverse tasks cannot reliably perform tasks of moderate length. Due to our limited dataset, we cannot confidently measure time horizons at very high success rates (e.g. 95%)– see Section 8.1.
@@ 6 External validity and robustness
To investigate the applicability of our results to other benchmarks, and to real task distributions, we performed four supplementary experiments. First, we check whether the 2023–2025 trend without the SWAA dataset retrodicts the trend since 2019, and find that the trends agree. Second, we label each of our tasks on 16 “messiness” factors—factors that we expect to (1) be representative of how real-world tasks may systematically differ from our tasks and (2) be relevant to AI agent performance. Third, we calculate AI agent horizon lengths from SWE-bench Verified tasks. We find a similar exponential trend, although with a shorter doubling time. However, we believe this shorter doubling time to be a result of SWE-bench Verified time annotations differentially underestimating the difficulty easier SWE-bench tasks. Finally, we collect and baseline a small set of uncontaminated issues from internal METR repositories. We find that our contracted human baseliners take much longer to complete these tasks than repository maintainers. We also find that AI agent performance is worse than would be predicted by maintainer time-to-complete but is consistent with contractor time-to-complete, given the AI agent success curves from HCAST + SWAA + RE-Bench tasks shown in Figure 5.
@@ 6.3 SWE-bench Verified
SWE-bench Verified time annotations were based on the expected time it would take an “engineer who has had a few hours to familiarize themselves with the codebase” to solve the issue [20]. We found that annotator time estimates differentially underestimate how long our contract baseliners take to complete the easiest SWE-bench verified tasks. As a result, our time horizon estimates for SWE-bench Verified (which use annotator times) are likely to underestimate the time horizon of less capable models relative to contractor times, in turn shortening doubling times. For more details see Appendix D.3.
@@ 7.2.2 Future changes in time horizon trends
The main inputs to AI research and development are compute and researcher time. If future AI systems are capable of substituting for human research engineers and/or increasing the compute-efficiency of training, the rate of AI progress will increase. We think it is likely that there will be substantial AI R&D automation once frontier AI time horizon reaches tens of hours, shortening the time from then until one-month-horizon AI.
@@ 8.1 Measuring and interpreting time horizon
An infinite time horizon does not mean an arbitrarily capable AI, merely the ability to complete tasks that take humans an arbitrarily long length of time. If an artificial general intelligence (AGI) is capable of completing all tasks expert humans can with a success rate of at least X%, its Xtime horizon will necessarily be infinite. Therefore, if such systems are ever developed, the long-term trend in time horizon will be faster than exponential, with an asymptote at the date of AGI deployment.
@@ 8.2 Limitations and future work
Data analysis
The 2024–2025 trend appears faster than the 2019–2025 trend, but due to the small number of models it is unclear if this is noise. Future work should include hypothesis testing to detect a possible 2024 slope change, and more sophisticated statistical methods to create credible intervals for an overall forecast. We also lose some information in the multiple stages of estimation—conversion of baseline data to task difficulty ratings, time horizon computation, and linear regression to find the trend—and end-to-end methods could be more data-efficient.
@@ Acknowledgments
The authors thank the following reviewers for feedback on draft versions of this paper. Ryan Greenblatt, Aaron Scher, Romeo Dean, Mike Knoop, Jeff Wu, Steve Newman, Rohit Krishnan, Taren Stinebrickner-Kauffman, JS Denain, Jacob Pfau, Seb Krier, Anton Troynikov, Max Henderson, Ajeya Cotra, Max Nadeau, Tamay Besiroglu, Nate Thomas. The authors thank Charles Foster and Michael Chen for their support with the publication.
@@ Appendix D More ablations and robustness checks
    Different weightings of task families: 1family $\frac{1}{\sqrt{\text{family size}}}$ vs uniform (included in Figure 12); also 1family $\frac{1}{\text{family size}}$
@@ D.1 Alternative curve fits
  • Sophisticated stochastic models that asymptote to infinity often have large uncertainties; see Roodman [2020], which applied a superexponential diffusion model to economic data. As we discuss in Section 7.2, large uncertainty may be appropriate when constructing a prediction interval for a long-term forecast, but quantitatively modeling the potential impact of the factors we mentioned is out of scope of this paper, so we prefer to discuss these qualitatively.
@@ D.3 SWE-bench Verified
SWE-bench Verified contains time estimates created by contractors that split tasks into four buckets: $<15$ min fix, 15 minutes–1 hour, 1 hour–4 hours, or $>4$ hours. These estimates were based on the expected time it would take an “engineer who has had a few hours to familiarize themselves with the codebase” to solve the issue.
We verified annotator time buckets by running seven baselines across 6 SWE-bench Verified tasks. We baselined four tasks in the “$<15$ min fix” bucket, and found that they took our baseliners 8, 26, 67, and 84 minutes respectively. We baselined two tasks in the “1 hour–4 hours” bucket, and both baseliners took between 2–3 hours.
@@
$>4$ hours 480.0 min
Table 8: We convert the SWE-bench Verified time annotations into task estimates, by taking the geometric mean of the time annotation. We caution that this likely underestimates the time each issue takes to resolve by a human baseliner without context—notably, we observe that the geometric mean of baseliner time for four randomly sampled tasks in the “$<15$ minute fix” bucket is 32.9 minutes.
@@
As shown in Table 8, we found that annotator time estimates are likely more inaccurate on the “$<$15 min fix” time bucket, as in practice baseliners often take much longer than this to complete tasks. This results in easier tasks’ difficulty being underestimated, meaning that the SWE-bench Verified time horizon estimates likely underestimate true model performance for less capable models, which in turn leads to a steeper exponential trend.
@@ Appendix F Author contributions
Brian Goodrich contributed to the development of the modular-public agent, and created the flock and duet agents. Nikola Jurkovic baselined SWAA and RE-Bench tasks and tested RE-Bench tasks. Seraphina Nix contributed to writing and revising the paper and figures. David Rein managed the development and collection of HCAST tasks and oversaw the human baselining process for those tasks. Lucas Jun Koba Sato led the development of the AI agents and collection of agent results on HCAST tasks. Chris Painter suggested substantial framing and presentation changes. Neev Parikh contributed to running external validity experiments.